OpenStudy (anonymous):

Given f(x) = –5x + 1, what is f(–2)?

jimthompson5910 (jim_thompson5910):

Everywhere you see an x, replace it with -2 and evaluate

jimthompson5910 (jim_thompson5910):

f(x) = –5x + 1 f(-2) = –5(-2) + 1 f(-2) = ???

OpenStudy (anonymous):

f(-2) = 11?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

One more? What is the slope of the graph of 2x – 6y = 9? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you able to isolate y?

OpenStudy (anonymous):

No...

jimthompson5910 (jim_thompson5910):

The first step is to move that 2x over what can we do to both sides to get rid of it on the left side?

OpenStudy (anonymous):

Subtract 9?

jimthompson5910 (jim_thompson5910):

we want to get rid of the 2x on the left side

jimthompson5910 (jim_thompson5910):

basically we want to replace 2x with 0 (so it goes away)

jimthompson5910 (jim_thompson5910):